Simplify. Write your answer with positive exponents only.
(b^2)^-6/b^-4
Please help me with this problem.
Thanks

Answer by stanbon(75887)

Write your answer with positive exponents only.
(b^2)^-6/b^-4
---------------
The negative in the exponent means "invert".
= b^4/(b^2)^6
= b^4/b^12
= 1/b^8
==============
Cheers,
Stan H.